A small block of mass 100 g moves with uniform speed in a horizontal circular groove, with vertical side walls, of radius 25 cm. If the block takes 2 s to complete one round, find the normal contact force by side wall of the groove.

Grade:12th pass

2 Answers

Apoorva Arora IIT Roorkee
9 years ago
The speed of the block is

v = \frac{2\Pi(25 cm) }{2 s} = 0.785 m/s.

The acceleration of the block is

a = \frac{v^{2}}{r} = \frac{0.785}{0.25} = 2.5 m/s^{2}

towards the centre. The only force in this direction is the normal contact force due to the side walls. Thus, from Newton’s second law, this force is

N = ma = (0.100)(2.5 m/s^{2}) = 0.25 N
Saileshwar
4 years ago

X+Y = a + b and X is considered asπ and y as ? so,we can add this two value and find this answer.